On Symmetric Block Designs (40,13,4) with Automorphisms of Order 13 Vladimir Ćepulić, Zagreb, Croatia

Keywords: symmetric block design, automorphism group, orbital strusture, indexing AMS Subject classification: 05B05 Abstract: All symmetric block designs (40,13,4) admitting an automorphism group of order 13 are classified. Also the generators and orders of their automorphism groups are determined.